# Modelling the underlying event in photon-initiated processes

### v1

@@ 1 Introduction
All the above models and implementations have been widely used in proton-proton collisions at the LHC, having in most cases been previously tuned to data from the Tevatron [7]. The case of photon-induced processes was studied for the Herwig and Pythia models several years ago [8] using HERA, LEP and TRISTAN data, and more recently the Sherpa implementation has been confronted with data from HERA and LEP [4]. Here we perform a similar study, with a somewhat wider range of data, for the current Pythia implementation. As with the Sherpa study, a prime motivation is the relevance of these models to upcoming data from the Electron Ion Collider [9]. We also note their relevance to photon-induced collisions at the LHC in both proton and heavy ion running [10, 11, 12].
@@ 2 The PYTHIA Model
The photon may interact either directly, or after developing a hadronic structure which is resolved in the collision. In the direct contribution, the photon scatters with a parton from the proton beam. In the resolved photon case, one needs to account for the partonic structure of (almost) real photons and sample a parton from the photon PDFs that participates in the hard process. In this case, the partonic evolution of the photon-proton subsystem proceeds as any hadronic collision including generation of multiparton interactions (MPIs) and parton-shower emissions. There are, however, a few differences that should be accounted for. First of all, the DGLAP evolution equation for resolved photons includes a so-called anomalous contribution that describes the perturbative splittings of a photon to quark-antiquark pair. To remain consistent with the resolved-photon PDFs, a similar term has been included into initial-state radiation algorithm within the default parton shower in Pythia. Having this term in the backwards evolution allows a resolved-photon state to collapse back to an unresolved state during the partonic evolution. The parton showers are generated simultaneously with the MPIs and further interactions below the scale at which the photon has become unresolved are rejected. If the photon remains resolved until the minimum parton-shower emission scale, remnant partons are added to conserve colour and momentum.
@@
The rate of MPIs also depends on the overlap of the colliding particles in the impact-parameter space [14]. In the case of collisions with photons, one could expect a different distribution of the spatial overlap compared to a proton-proton collisions. Here we, however, consider only the variation of $p_{\mathrm{T,0}}$, and its energy dependence, that effectively accounts also for the modified spatial structure of the collision in the context of MPI generation. In addition, this spatial overlap is also affected by the hardness of primary scattering as discussed in the original article [1].
@@
This is the default parametrization in Pythia for photon-photon collisions [16] based on a tune to LEP data for charged-particle production in $\gamma\gamma$ collisions [17] at $10<W<125~{}\text{GeV}$.
@@ 3.1 Definition of Kinematic Variables
The mean pseudorapidity of the two jets, $\bar{\eta}$, is given by
@@ 4 Jet Production Data
A comparison to the lowest-transverse-energy jet data, available from ZEUS and OPAL, is shown in Fig. 2. It can be seen that while MPIs are required to describe the data, and the LEP and LHC tunes with logarithmic energy scaling both do a reasonable job, the Detroit and 2C models, tuned to describe proton-antiproton collisions at energies closer to the average photon-proton centre-of-mass energy at HERA and the average photon-photon energy at LEP2, do not model the data well; there is too much additional transverse energy in the events, enhancing the jet cross sections to the extent that they are far above the data, especially in regions where resolved photon interactions dominate – that is, at high rapidity at HERA (due to the asymmetric kinematics) and at low $x_{\gamma}^{\rm obs}$ at LEP2.
Figure 2: (Left) HERA photoproduction dijet data ($\gamma p$1) [23], where both jets are required to have transverse energies greater than 6 GeV, and (right) LEP dijet data ($\gamma\gamma$1) [29] compared to models of the underlying event in Pythia tuned to $\gamma\gamma$ data (LEP/LOG), LHC $pp$ data with modified energy dependence (LHC/LOG) and low-energy $pp$ data (Detroit, 2C), and to results without MPIs.
@@ 5 Charged Hadron Production
We now move away from jet production to the charged particle production which is potentially more sensitive to the MPI tuning since these give access to lower values of $p_{\mathrm{T}}$, though at the price of a less direct connection to hard-scattering kinematics.
Attempts have previously been made to tune the value of $p_{\mathrm{T},0}^{\rm ref}$ to HERA [28] and LEP data [17]. A comparison with LEP data, a study [34] of charged hadron production yielded a value of about 3.3 GeV and in comparison to HERA data, a value of about 3.0 GeV was found to best describe the data. A complementary study using newer ZEUS inclusive charged particle data [35] also indicates that a value of about 3.0 GeV is preferred. These values are above the default used for LHC data of 2.28 GeV and indicate that fewer MPIs are present in photon-initiated processes, as was observed in the case of dijet production above.
@@
Finally we show these tunes compared to charged-hadron production in the underlying event at the LHC, for three different centre-of-mass collision energies – 900 GeV, 7 TeV [30], and 13 TeV [31] in Fig. 9. While none of the parameter settings describe the data perfectly the two LHC tunes are closer than the LEP tunes, which lie well below the data at all collision energies. We notice that the LHC/POWER tune retains approximately the same level of agreement with the data at all three collision energies, which is expected, since the tune made use of the lower-energy data as a constraint for the MPI parameters. The LHC/LOG tune moves from being below the data at 900 GeV to being above it at 13 TeV. This suggests that a power-law energy dependence is appropriate in proton-proton collisions. However, we have shown that this parameterization (LHC/POWER) gives poor agreement with the jet data from photon-induced collisions.
@@ 6 Discussion, Recommendations and Summary
Describing jet and charged-particle production in high-energy colliders, especially at relatively low transverse momenta, is challenging for the Monte Carlo simulations commonly used in experimental design, in measurements and in interpretation of the data. Our study confirms that multiparton interaction models are an essential component of any reasonable description of these data, also in the case of processes initiated by low-virtuality photons.
@@
So far, tuning efforts for the new implementations of photoproduction and $\gamma\gamma$ collisions in MC event generators have been sparse. In the Pythia context, a default set of parameters for $\gamma\gamma$ collisions were derived based on LEP charged-particle production data that can also fairly well describe the LEP jet production data, as shown in this study. However, this study also shows that applying this parametrization to HERA photoproduction data leads to an underestimate the cross sections for both charged-particle and dijet production. This becomes especially evident in the case of multijet data from ZEUS, which seem to be very sensitive to the underlying MPI tune.
@@
In addition to MPI tuning, there are potentially other modelling aspects that could improve the description of the data with the current Pythia implementation. For example we notice that HERA $x_{\gamma}^{\rm obs}$ measurements in the $0.80.6<x_{\gamma}^{\rm obs}<$ 0.8 region and for high-$E_{\textrm{T}}$ jets are poorly described in all parameter sets. Potentially this could be improved by re-fitting real-photon PDFs using more recent data and theory, especially from HERA. Similarly, increasing the perturbative precision in the hard-process matrix elements and in the parton showers could potentially improve the description of the data in the regions where non-MPI related discrepancies occur.

### v2

@@ 1 Introduction
All the above models and implementations have been widely used in proton-proton collisions at the LHC, having in most cases been previously tuned to data from the Tevatron [7]. The case of photon-induced processes was studied for the Herwig and Pythia models several years ago [8] using HERA, LEP and TRISTAN data, and more recently the Sherpa implementation has been confronted with data from HERA and LEP [4]. Here we perform a similar study, with a somewhat wider range of data, for the current Pythia implementation. As with the Sherpa study, a prime motivation is the relevance of these models to upcoming data from the Electron Ion Collider [9]. We also note their relevance to photon-induced collisions at the LHC in both proton and heavy-ion running [10, 11, 12].
@@ 2 The PYTHIA Model
The photon may interact either directly, or after developing a hadronic structure which is resolved in the collision. In the direct contribution, the photon scatters with a parton from the proton beam. In the resolved photon case, one needs to account for the partonic structure of (almost) real photons and sample a parton from the photon PDFs that participates in the hard process. In this case, the partonic evolution of the photon-proton subsystem proceeds as any hadronic collision including generation of multiparton interactions (MPIs) and parton-shower emissions. The possibility of MPIs is required to describe the data [14]. There are, however, a few differences that should be accounted for. First of all, the DGLAP evolution equation for resolved photons includes a so-called anomalous contribution that describes the perturbative splittings of a photon to quark-antiquark pair. To remain consistent with the resolved-photon PDFs, a similar term has been included into initial-state radiation algorithm within the default parton shower in Pythia. Having this term in the backwards evolution allows a resolved-photon state to collapse back to an unresolved state during the partonic evolution. The parton showers are generated simultaneously with the MPIs and further interactions below the scale at which the photon has become unresolved are rejected. If the photon remains resolved until the minimum parton-shower emission scale, remnant partons are added to conserve colour and momentum.
@@
The rate of MPIs also depends on the overlap of the colliding particles in the impact-parameter space [15]. In the case of collisions with photons, one could expect a different distribution of the spatial overlap compared to proton-proton collisions. Here we, however, consider only the variation of $p_{\mathrm{T,0}}$, and its energy dependence, that effectively accounts also for the modified spatial structure of the collision in the context of MPI generation. In addition, this spatial overlap is also affected by the hardness of primary scattering as discussed in the original article [1].
@@
This is the default parametrization in Pythia for photon-photon collisions [17] based on a tune to LEP data for charged particle production in $\gamma\gamma$ collisions [18] at $10<W<125~{}\text{GeV}$.
@@ 3.1 Definition of Kinematic Variables
The mean pseudorapidity of the two jets in the laboratory frame, $\bar{\eta}$, is given by
@@ 4 Jet Production Data
A comparison to the lowest-transverse-energy jet data, available from ZEUS and OPAL, is shown in Fig. 2. It can be seen that while MPIs are required to describe the data, as mentioned previously, and the LEP and LHC tunes with logarithmic energy scaling both do a reasonable job, the Detroit and 2C models, tuned to describe proton-antiproton collisions at energies closer to the average photon-proton centre-of-mass energy at HERA and the average photon-photon energy at LEP2, do not model the data well; there is too much additional transverse energy in the events, enhancing the jet cross sections to the extent that they are far above the data, especially in regions where resolved photon interactions dominate – that is, at high rapidity at HERA (due to the asymmetric kinematics) and at low $x_{\gamma}^{\rm obs}$ at LEP2.
Figure 2: (Left) HERA photoproduction dijet data ($\gamma p$1) [24], where both jets are required to have transverse energies greater than 6 GeV, and (right) LEP dijet data ($\gamma\gamma$1) [29] compared to models of the underlying event in Pythia tuned to $\gamma\gamma$ data (LEP/LOG), LHC $pp$ data with modified energy dependence (LHC/LOG) and low-energy $pp$ data (Detroit, 2C), and to results without MPIs. The error bars on the Pythia expectations shown here and in subsequent figures represent statistical uncertainties.
@@ 5 Charged Hadron Production
We now move away from jet production to charged particle production which is potentially more sensitive to the MPI tuning since these give access to lower values of $p_{\mathrm{T}}$, though at the price of a less direct connection to hard-scattering kinematics.
Attempts have previously been made to tune the value of $p_{\mathrm{T},0}^{\rm ref}$ to HERA [28] and LEP data [18]. In comparison with LEP data, a study [34] of charged hadron production yielded a value of about 3.3 GeV and in comparison to HERA data, a value of about 3.0 GeV was found to best describe the data. A complementary study using newer ZEUS inclusive charged particle data [35] also indicates that a value of about 3.0 GeV is preferred. These values are above the default used for LHC data of 2.28 GeV and indicate that fewer MPIs are present in photon-initiated processes, as was observed in the case of dijet production above.
@@
Finally we show these tunes compared to charged hadron production in the underlying event at the LHC, for three different centre-of-mass collision energies – 900 GeV, 7 TeV [30], and 13 TeV [31] in Fig. 9. While none of the parameter settings describe the data perfectly the two LHC tunes are closer than the LEP tunes, which lie well below the data at all collision energies. We notice that the LHC/POWER tune retains approximately the same level of agreement with the data at all three collision energies, which is expected, since the tune made use of the lower-energy data as a constraint for the MPI parameters. The LHC/LOG tune moves from being below the data at 900 GeV to being above it at 13 TeV. This suggests that a power-law energy dependence is appropriate in proton-proton collisions. However, we have shown that this parameterization (LHC/POWER) gives poor agreement with the jet data from photon-induced collisions.
@@ 6 Discussion, Recommendations and Summary
Describing jet and charged particle production in high-energy colliders, especially at relatively low transverse momenta, is challenging for the Monte Carlo simulations commonly used in experimental design, in measurements and in interpretation of the data. Our study confirms that multiparton interaction models are an essential component of any reasonable description of these data, also in the case of processes initiated by low-virtuality photons.
@@
So far, tuning efforts for the new implementations of photoproduction and $\gamma\gamma$ collisions in MC event generators have been sparse. In the Pythia context, a default set of parameters for $\gamma\gamma$ collisions were derived based on LEP charged particle production data that can also fairly well describe the LEP jet production data, as shown in this study. However, this study also shows that applying this parametrization to HERA photoproduction data leads to an underestimate the cross sections for both charged particle and dijet production. This becomes especially evident in the case of multijet data from ZEUS, which seem to be very sensitive to the underlying MPI tune.
@@
In addition to MPI tuning, there are potentially other modelling aspects that could improve the description of the data with the current Pythia implementation. For example we notice that HERA $x_{\gamma}^{\rm obs}$ measurements in the $0.80.6<x_{\gamma}^{\rm obs}<$ 0.8 region and for high-$E_{\textrm{T}}$ jets are poorly described in all parameter sets. Potentially this could be improved by re-fitting real-photon PDFs using more recent data and theory, especially from HERA. Explicitly including the possibility of a direct photon interaction with multiple partons may also have an impact [37]. Similarly, increasing the perturbative precision in the hard-process matrix elements and in the parton showers could potentially improve the description of the data in the regions where non-MPI related discrepancies occur.